OpenStudy (anonymous):

how many solutions does the system of equations have 2x = -10y + 6 and x + 5y = 3 one two infinitely many none

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

2x = -10y + 6 2x + 10y = 6 reduces to x + 5y = 3 same line as other one, therefore, infinite many solutions

OpenStudy (anonymous):

can you help me with two more questions please?

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

y - 6x = -3 and 4y - 24x = -16 one two infinitely many none

OpenStudy (texaschic101):

4y - 24x = -16 reduces to y - 6x = -4 so we have : y - 6x = -3 and y - 6x = -4 there is no solutions to this one

OpenStudy (anonymous):

ok im going to post a graph row quick

OpenStudy (anonymous):

what is the graph of the inequality in the coordinate plane x ≥ 3

OpenStudy (anonymous):

I think its two

OpenStudy (texaschic101):

x >= 3....this is a vertical line. and it is a vertical line greater then 3, so you are correct, graph 2

OpenStudy (anonymous):

yay :)

OpenStudy (texaschic101):

you deserve a medal for that :)

OpenStudy (anonymous):

I have two more how many solutions questions do you wanna help me with them?

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

y = 4x + 3 and 2y - 8x = 3 one two infinitely many none

OpenStudy (texaschic101):

y = 4x + 3 -4x + y = 3 4x - y = -3 -8x + 2y = 3 8x - 2y = -3 4x - y = -3 -->(-2) 8x - 2y = -3 -------------- -8x + 2y = 6 (result of multiplying by -2) 8x - 2y = -3 ------------add 0 = 3 (incorrect) I do not think there is a solution to this one

OpenStudy (anonymous):

looking at that I don't think theres a solution either

OpenStudy (anonymous):

last one :) x - 4y = 12 and 5x - 20y = 60

OpenStudy (texaschic101):

I had to work on that one a little while..lol

OpenStudy (anonymous):

same possible answers one two infinitely many none

OpenStudy (texaschic101):

5x - 20y = 60 reduces to x - 4y = 12 same line...infinite solutions

OpenStudy (anonymous):

thank you so much! you are a lifesaver

OpenStudy (texaschic101):

glad I could help :)

OpenStudy (anonymous):

:) thank you

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

Bye! Ill message you later if I need help :))))